OpenStudy (anonymous):

In his will, Uncle Clarence left $5000 to his two nieces. Emma’s share is to be $500 more than Clara’s. How much should each niece receive?

OpenStudy (anonymous):

Your total is going to be set equal to 5000. Let x be the amount Clara gets, and x+500 is the amount Emma gets. Set up the equation with that information.

OpenStudy (anonymous):

5000-x=500+x +x +x 5000=500+2x -500 -500 \[\frac{ 4500 }{ 2 } = \frac{ 2x }{ 2 }\] x=2250 Clara's share $2,250 Emma's share $5,000 - 2250 = $2,750 Is this right?